The masses of the Earth and the moon are 5.97 x 1024 kg and 7.36 x 1022 kg, respectively, while the average Earth-Moon distance is 3.84 x108 m.


What gravitational force exist between them?

1
Expert's answer
2021-11-01T12:53:19-0400

Gravitational force for Earth and Moon system:

"F_G = \frac{GMm}{R^2}"

G = gravitational constant

M = mass of Earth

m = mass of Moon

R = distance between them

"G = 6.674 \times 10^{-11} \\

M = 5.97 \times 10^{24} \;kg \\

m = 7.36 \times 10^{22} \;kg \\

R = 3.84 \times 10^8 \;m \\

F_G = \frac{6.674 \times 10^{-11}\times 5.97 \times 10^{24} \times 7.36 \times 10^{22}}{(3.84 \times 10^8)^2} \\

F_G = 1.986 \times 10^{20} \;N"